OpenStudy (anonymous):

A student scored 78 and 95 on his first two quizzes. Use a compound inequality to find the possible values for a third quiz score that would give him an average between 80 and 90, inclusive.

OpenStudy (anonymous):

67 <= x <= 97 78 <= x <= 95 80 <= x <= 90 86 <= x <= 87

OpenStudy (anonymous):

Multiply the desired average by 3 for the 3 tests then subtract the two known values.

OpenStudy (anonymous):

I'm sorry I'm really lost.

OpenStudy (anonymous):

To get an average of 80: \(80 \times 3 = 240\) So you need a total of 240 points. \(240 - 78=162; 162-95=67\) Therefore, you would need a 67 on the 3rd test to get the lower average of 80.